# A Survey on Data Selection for Language Models

### v2

@@ 1 Introduction
Data selection is a long-standing challenge of machine learning where, given a collection of raw data, the goal is to design a dataset that is optimal in some sense (John & Draper, 1975).
@@
One often used sense of optimality in data selection is with regard to a model’s performance. In this work, we adopt the commonly held view that, at their core, machine learning models are a method for modeling statistical patterns in data and, from the probabilistic viewpoint, the optimal dataset is that which most closely matches the distribution under which the model will be evaluated (Murphy, 2012). While the probabilistic viewpoint presents a common view of how to select data that improves model performance, this is not the only goal of data selection methods. Data selection methods can reduce costs (by reducing dataset size (Ortiz Suárez et al., 2019; Schreiber et al., 2020; Brown et al., 2020; Lee et al., 2022a; Sorscher et al., 2022)), ensure the integrity of evaluation metrics (by removing data that is suspected to be from the evaluation data (Rae et al., 2021; Marone & Van Durme, 2024; Oren et al., 2024)), and reducing undesirable behaviors (such as bias and toxicity (Dodge et al., 2021; Welbl et al., 2021; Luccioni & Viviano, 2021; Longpre et al., 2023c)).
@@
Data selection has recently become especially important in the context of large language models. Language models can undergo multiple stages of training (pretraining (Peters et al., 2018; Radford & Narasimhan, 2018; Devlin et al., 2019; Raffel et al., 2020; Touvron et al., 2023a), instruction-tuning (Mishra et al., 2021; Sanh et al., 2022; Longpre et al., 2023a; Muennighoff et al., 2024), alignment (Ziegler et al., 2019; Bai et al., 2022b; Ouyang et al., 2022; Rafailov et al., 2023), etc.), and data selection plays an important role in each stage. However, as the objectives of training differ across each stage, the goals of data selection also vary accordingly. For example, language models are commonly pretrained on large corpora of text from a variety of sources. Among these sources is the internet, where estimates suggest that there are around 250 billion webpages. These webpages amount to around 11 petabytes of data, collected from internet scraping efforts since 2008, with an additional 3-5 billion new web pages being crawled monthly. [^1] According to https://commoncrawl.org/ accessed on 11/02/23 (MM/DD/YY). Due to the massive size of pretraining corpora, a common goal of data selection during pretraining is to remove significant quantities of data through a series of filters (Conneau & Lample, 2019; Raffel et al., 2020; Wenzek et al., 2020; Gao et al., 2020; Rae et al., 2021; Lee et al., 2022a) that aim to only retain data that is deemed “high-quality”. In contrast with pretraining, one form of data selection for fine-tuning a model on a target task is to select additional auxiliary samples that will be most beneficial as additional learning signals for the target task (Albalak et al., 2023b; Ivison et al., 2023a).
@@
In this work, we unify the wide array of data selection methods under a conceptual framework that allows us to compare and contrast the variety of methods under our probabilistic viewpoint (§2.2), with a focus on model pretraining. Through this survey, we demonstrate how all data selection methods define a utility function that determines the utility of data, as well as a selection mechanism that determines how to use a data point based on its utility. Our conceptual framework enables us to classify the wide variety of data selection methods and create a taxonomy of the existing literature. Overall, this survey aims to achieve two goals: 1) we provide a collected resource of data selection methods that describes the current best practices and considerations when selecting data for training a language model, and 2) we provide a unifying view of data selection methods that allows us to define and describe potentially fruitful future directions of research. While this survey aims to be comprehensive, it would be prohibitively long to include the exact details of every method, so we select a representative sample to discuss in depth and provide citations to the many methods that we cannot cover in depth.
@@
We organize the survey as follows. First, we present a taxonomy and unified conceptual framework for data selection (§2). Next, we present the main focus of this work: surveying methods of data selection for language model pretraining (§3). Then, we follow up with data selection methods for other language model training regimes including multitask training and instruction-tuning (§4), alignment (§5), in-context learning (§6), and task-specific fine-tuning (§7). Next, we extend the survey to notable methods of data selection used in domains other than language (§8), as well as a brief discussion of related topics(§9). Then, we discuss some of the implications surrounding data selection (§10). Although the main focus of the survey is on language model pretraining, we also cover methods from other training regimes and domains so that we can point out promising directions of research in the final section (§11).
@@
Figure 1: An overview of the data pipeline for language models. The process starts with raw data, that is cleaned, filtered, and mixed to create a final dataset by the data selection process, then used to train (or evaluate) a model. The details and objectives of data selection methods vary depending on the learning stage, and we identify five common objectives: improving model performance, improving data efficiency, selecting data efficiently, ensuring evaluation integrity, and reducing model bias and toxicity. For example, when selecting data for pretraining (§3) we may care about data and selection efficiency more than about reducing toxicity in the data. In addition to pretraining, data selection can be used for instruction-tuning (§4), alignment (§5), in-context learning (§6), and task-specific fine-tuning (§7), where each stage has differing priorities.
@@ 2.1 Background and Motivation of Data Selection
Prior to discussing data selection methods, we first define and describe the “units” of a dataset, from smallest to largest. First, the smallest unit of data for language models is the token, which can be composed of bytes, characters, subwords, or larger components, depending on the tokenizer.
@@
A data point, $x^{(i)}$, is an ordered collection of tokens that constitutes a single sample of data used to train or evaluate a model.
For example, in language modeling, $x^{(i)}$ can be a sequence of tokens from an internet document, a book, or a scientific article. Language models have a limited input sequence length when training, so long documents are often split into multiple data points. In practice, this means that $x^{(i)}$ may constitute a portion of a document (rather than a complete one).
@@
The data point characteristics are a crucial component of data selection as they are frequently used to determine whether a data point needs to be cleaned or possibly removed entirely. A number of different measures can be used to describe individual data points. Common measures used to describe data points can be either simple statistics (e.g. number of characters in $x^{(i)}$, or the percent of alphabetic characters in $x^{(i)}$) or a distributed representation (e.g. the embedded representation using BERT (Devlin et al., 2019)).
@@
The distribution of a dataset, $\mathcal{D}$, refers to the distribution of data points within the space of all data possible data points. The distribution of $\mathcal{D}$ has a significant impact on the final capabilities and properties of a model that is trained on it.
The dataset distribution is also a crucial component of data selection as models can struggle with out-of-distribution generalization and the dataset distribution specifies what data is in-distribution vs. out-of-distribution (Shi et al., 2023). Specifically, a dense region of the data distribution generally suggests that a model will perform well on unseen data from that region (e.g. similar distribution of tokens). With this in mind, increasing the density of data around desirable regions while reducing density around undesirable regions generally leads to a model that performs well in the desired settings.
@@ 2.2 A Unified Conceptual Framework for Data Selection
Data selection is the process of creating a dataset from a collection of candidate data points, which will be used to train or evaluate a machine learning model. Prior to data selection, generally speaking, data is collected, annotated, and stored. Then, once the target use case for a model is determined, the goal of data selection is to filter and select the data points that maximize the desired objective.
@@
A data selection function, $\phi$, takes as input a dataset $\mathcal{D}_{\mathrm{raw}}$ and objective function $f_{\mathrm{obj}}$ and filters, cleans, and selects data points from $\mathcal{D}_{\mathrm{raw}}$ to create a final dataset $\phi(\mathcal{D}_{\mathrm{raw}})=\mathcal{D}$ such that, for a model $\mathcal{M}$ trained on dataset $\mathcal{D}$, $\phi$ aims to maximize/minimize the objective function $f_{\mathrm{obj}}(\mathcal{M})$.
@@
The selection mechanism uses the output from the utility function to determine whether a data point will be included in the resulting subset (and in some cases, how many times the data point should be repeated). It can be a simple indicator (e.g., only include $x^{(i)}$ in $\mathcal{D}$ if the number of characters is greater than 10), or probability functions (e.g. if $\upsilon$ is the likelihood of $x^{(i)}$ being a Wikipedia article, include $x^{(i)}$ in $\mathcal{D}$ according to the probability defined by $\upsilon(x^{(i)}$)). Additionally, a selection sensitivity is needed for selection mechanisms that require a threshold (e.g. only include $x^{(i)}$ if $\upsilon(x^{(i)})>0.9$).
@@ 2.3 Dimensions of Variance in the Data Selection Taxonomy
The main goal of distribution matching methods is to select data with properties similar to the desired target distribution, upon which a model will be evaluated or deployed. For instance, the desired distribution could be defined as data of known high quality, a specific language, or a target domain (e.g finance, medicine, or law). The exact specifications of the desired target distribution can vary from being well defined (e.g. as in detecting a language) to being quite ambiguous (e.g. “high quality” data). Some distribution matching methods will try to match data representation distributions where the similarity to the target distribution will usually be the utility (e.g. similarity to Wikipedia data). Other distribution matching methods use the statistics of data sampled from the target dataset as the utility (e.g. total number of characters per example).
Distribution diversification methods aim to prioritize heterogeneity in a sample, removing redundancies. They operate within a representation space that allows for similarity to be measured across data points. The utility of a data point is defined by its relation (similarity) to the other data points in the representation space. Distribution diversification methods often remove data points that are similar in some representation space (e.g. characters or vectors). By removing data points that are very similar, diversification methods can remove redundancies in the data, leading to improved training efficiency by reducing the dataset size. Additionally, because distribution diversification methods force the dataset distribution to be flatter, they can lead to decreased memorization, decreased bias, and improved robustness.
@@
Methods that alter the dataset aim to increase or decrease the frequency of individual data points within the dataset in order to increase (decrease) the resultant distribution density around desirable (undesirable) regions. These methods take as input a dataset and assign each individual data point a non-negative integer value representing the number of times that data point should be included in the resultant dataset. Formally, a dataset distribution altering function $\phi_{j}:\mathcal{D}^{N}\to\mathbb{N}_{0}^{N}$ maps each data point $x^{(i)}\in\mathcal{D}$ to a non-negative integer based on the utility function, $\upsilon(x^{(i)})$.
Methods that alter the data point aim to adjust the content within a data point to better match a desirable distribution of tokens so that the individual data points within a dataset more closely resemble the desired characteristics. In practice, this can take the shape of removing individual lines, or chunks of text, from a document. For example, when training a model for natural language, it can be beneficial to remove lines that contain HTML tags as they are likely to be out of distribution for the desired target domain.
@@
Methods that alter the dataset distribution lead to another dimension of variance; whether the selection function $\phi$ assigns binary values (i.e. include or remove) or whether larger integer values are permitted. Methods that only assign binary values are generally referred to as filtering, and the goal is usually to adjust the dataset distribution by removing undesirable data points from the dataset. On the other hand, methods that can assign any natural number are commonly referred to as data mixing, and their goal is often to adjust the dataset distribution by prioritizing certain subsets of the dataset that have high utility values, while decreasing the distribution density around subsets with lower utility. It is very common for data mixing methods to define a utility function that assigns the same utility to entire subsets of data (e.g. all web text gets the same utility, and all books get a different utility). Filtering and mixing can be used as part of a pipeline, where data is first filtered and then mixed.
@@
2.3.4 Training Stage
Each training stage has different goals, and so the data selection methods for each stage will use different mechanisms to achieve those goals. For example, the target distribution is fairly well defined when training a language model for a specific domain or task, but much less defined when pretraining a general purpose language model, thus data selection methods can utilize this information to target better data. Another consideration is the number of candidate data points at each training stage. For example, pretraining likely has a significantly larger number of candidate data points than instruction-tuning, motivating the need for efficient data selection methods in pretraining, while instruction-tuning can afford more expensive methods. In this work, we focus on data selection for pretraining.
3 Data Selection for Pretraining
@@
Figure 4: An overview of the data filtering pipeline for pretraining. Each filtering component is described in Section 3, and depicts common filters used for preprocessing text data. Note that different works employ different filters, at different stages, and do not necessarily adhere to the order conveyed here. This figure was adopted and modified from Soldaini et al. (2024)
The goal of pretraining is usually to train a general-purpose model, which requires training on massive quantities of text, commonly measured in the billions and trillions of tokens. Selecting the best data from such large quantities can be very expensive, so a common first step in the process is to remove data with various filters, and it is likely that multiple filters will need to be pipelined together to achieve the desired dataset. The order in which we present pretraining data selection methods is, roughly, based on the order that they are used within a real data selection pipeline. Of course, not all pipelines require every method presented here, and depending on the case the exact ordering may differ slightly.

### v3

@@ Abstract
*Work completed while at UC Santa Barbara. Correspondence to alon_albalak@ucsb.edu
@@ 1 Introduction
Data selection is a long-standing challenge of machine learning where, given a collection of raw data, the goal is to design a dataset that is optimal under some objective function (John & Draper, 1975).
@@
One often-used sense of optimality in data selection is with regard to a model’s performance. In this work, we adopt the commonly held view that, at their core, machine learning models are a method for modeling statistical patterns in data and, from the probabilistic viewpoint, the optimal dataset is that which most closely matches the distribution under which the model will be evaluated (Murphy, 2012). While the probabilistic viewpoint presents a common view of how to select data that improves model performance, this is not the only goal of data selection methods. Data selection methods can reduce costs (by reducing dataset size (Ortiz Suárez et al., 2019; Schreiber et al., 2020; Brown et al., 2020a; Lee et al., 2022a; Sorscher et al., 2022)), ensure the integrity of evaluation metrics (by removing data that is suspected to be from the evaluation data (Rae et al., 2021; Marone & Van Durme, 2024; Oren et al., 2024)), and reducing undesirable behaviors (such as bias and toxicity (Dodge et al., 2021; Welbl et al., 2021; Luccioni & Viviano, 2021; Longpre et al., 2023c)).
@@
Data selection has recently become especially important in the context of large language models (Zhao et al., 2023b; Minaee et al., 2024). Language models can undergo multiple stages of training (pretraining (Peters et al., 2018; Radford & Narasimhan, 2018; Devlin et al., 2019; Raffel et al., 2020; Touvron et al., 2023a), instruction-tuning (Mishra et al., 2021; Sanh et al., 2022; Longpre et al., 2023a; Muennighoff et al., 2024), alignment (Ziegler et al., 2019; Bai et al., 2022b; Ouyang et al., 2022; Rafailov et al., 2023), etc.), and data selection plays an important role in each stage. However, as the objectives of training differ across each stage, the goals of data selection also vary accordingly. For example, language models are commonly pretrained on large corpora of text from a variety of sources. Among these sources is Common Crawl, a collection of around 250 billion webpages scraped from the internet. These webpages amount to around 11 petabytes of data, collected from internet scraping efforts since 2008, with an additional 3-5 billion new web pages being crawled monthly. [^1] According to https://commoncrawl.org/ accessed on 11/02/23 (MM/DD/YY). Due to the massive size of pretraining corpora, a common goal of data selection during pretraining is to remove significant quantities of data through a series of filters (Conneau & Lample, 2019; Raffel et al., 2020; Wenzek et al., 2020; Gao et al., 2020; Rae et al., 2021; Lee et al., 2022a) that aim to only retain data that is deemed “high-quality”. In contrast with pretraining, one form of data selection for fine-tuning a model on a target task is to select additional auxiliary samples that will be most beneficial as additional learning signals for the target task (Albalak et al., 2023b; Ivison et al., 2023a).
@@
In this work, we unify the wide array of data selection methods under a conceptual framework that allows us to compare and contrast the variety of methods under our probabilistic viewpoint (§2.2) with a focus on model pretraining. Through this survey, we demonstrate how all data selection methods define a utility function that determines the utility of data, as well as a selection mechanism that determines how to use a data point based on its utility. Our conceptual framework enables us to classify the wide variety of data selection methods and create a taxonomy of the existing literature. Overall, this survey aims to achieve two goals: 1) we provide a collected resource of data selection methods that describes the current best practices and considerations when selecting data for training a language model, and 2) we provide a unifying view of data selection methods that allows us to define and describe potentially fruitful future directions of research. While this survey aims to be comprehensive, it would be prohibitively long to include the exact details of every method, so we select a representative sample to discuss in depth and provide citations to the many methods that we cannot cover in depth.
@@
We organize the survey as follows. First, we present a taxonomy and unified conceptual framework for data selection (§2). Next, we present the main focus of this work: surveying methods of data selection for language model pretraining (§3). Then, we follow up with data selection methods for other language model training regimes including multitask training and instruction-tuning (§4), alignment (§5), in-context learning (§6), and task-specific fine-tuning (§7). We then extend the survey to notable methods of data selection used in domains other than language (§8), as well as a brief discussion of related topics (§9). Then, we discuss some of the implications surrounding data selection (§10). Although the main focus of the survey is on language model pretraining, we also cover methods from other training regimes and domains so that we can point out promising directions of research in the final section (§11).
@@
Figure 1: An overview of the data pipeline for language models. The process starts with raw data that is cleaned, filtered, and mixed to create a final dataset by the data selection process, then used to train (or evaluate) a model. The details and objectives of data selection methods vary depending on the learning stage, and we identify five common objectives: improving model performance, improving data efficiency, selecting data efficiently, ensuring evaluation integrity, and reducing model bias and toxicity. For example, when selecting data for pretraining (§3) we may care about data and selection efficiency more than about reducing toxicity in the data. In addition to pretraining, data selection can be used for instruction-tuning (§4), alignment (§5), in-context learning (§6), and task-specific fine-tuning (§7), where each stage has differing priorities.
@@ 2.1 Background and Motivation of Data Selection
Prior to discussing data selection methods, we first define and describe the “units” of a dataset, from smallest to largest. First, the smallest unit of data for language models is the token, which can be composed of bytes, characters, subwords, or larger components, depending on the tokenizer. [^2] How to effectively encode texts is still an active research area (Rust et al., 2022), but to date splitting text into tokens is the leading approach.
@@
A data point, $x^{(i)}$, is an ordered collection of tokens that constitutes a single sample of data used to train or evaluate a language model.
For example, in language modeling, $x^{(i)}$ can be a sequence of tokens from an internet document, a book, or a scientific article. Language models often have a limited input sequence length when training, so long documents are often split into multiple data points. In practice, this means that $x^{(i)}$ may constitute a portion of a document (rather than a complete one).
@@
Common measures used to describe data points can be either simple statistics (e.g., number of characters in $x^{(i)}$, or the percent of alphabetic characters in $x^{(i)}$) or a distributed representation (e.g., the embedded representation using BERT (Devlin et al., 2019)). The data point characteristics are a crucial component of data selection as they are frequently used to determine whether a data point should be cleaned or possibly removed entirely.
@@
The distribution of a dataset, $\mathcal{D}$, refers to the distribution of data points within the space of all possible data points. The distribution of $\mathcal{D}$ has a significant impact on the final capabilities and properties of a model that is trained on it.
The dataset distribution is also a crucial component of data selection as models can struggle with out-of-distribution generalization and the dataset distribution specifies what data is in-distribution vs. out-of-distribution (Shi et al., 2023). Specifically, a dense region of the data distribution generally suggests that a model will perform well on unseen data from that region (e.g., similar distribution of tokens). With this in mind, increasing the density of data around desirable regions while reducing density around undesirable regions generally leads to a model that performs well in the desired settings.
@@ 2.2 A Unified Conceptual Framework for Data Selection
Data selection is the process of creating a dataset from a collection of candidate data points, which will be used to train or evaluate a machine learning model. Prior to data selection, generally speaking, data is collected, (optionally) annotated, and stored. Then, once the target use case for a model is determined, the goal of data selection is to filter and select the data points that maximize a desired objective.
@@
A data selection function, $\phi$, takes as input a dataset $\mathcal{D}_{\mathrm{raw}}$ and objective function $f_{\mathrm{obj}}$ and filters, cleans, and selects data points from $\mathcal{D}_{\mathrm{raw}}$ to create a final dataset $\mathcal{D}=\phi(\mathcal{D}_{\mathrm{raw}})$ such that, for a model $\mathcal{M}$ trained on dataset $\mathcal{D}$, $\phi$ aims to maximize/minimize the objective function $f_{\mathrm{obj}}(\mathcal{M})$.
@@
The selection mechanism uses the output from the utility function to determine whether a data point will be included in the resulting subset (and in some cases, how many times the data point should be repeated). It can be a simple indicator (e.g., only include $x^{(i)}$ in $\mathcal{D}$ if the number of characters is greater than 10), or probability functions (e.g., if $\upsilon$ is the likelihood of $x^{(i)}$ being a Wikipedia article, include $x^{(i)}$ in $\mathcal{D}$ according to the probability defined by $\upsilon(x^{(i)}$)). Additionally, a selection sensitivity is needed for selection mechanisms that require a threshold (e.g., only include $x^{(i)}$ if $\upsilon(x^{(i)})>0.9$).
@@ 2.3 Dimensions of Variance in the Data Selection Taxonomy
The main goal of distribution matching (Figure 2, left) methods is to select data with properties similar to the desired target distribution, upon which a model will be evaluated or deployed. For instance, the desired distribution could be defined as data of known high quality, a specific language, or a target domain (e.g., finance, medicine, or law). The exact specifications of the desired target distribution can vary from being well-defined (e.g., as in detecting a language) to being quite ambiguous (e.g., “high-quality” data). Some distribution matching methods will try to match data representation distributions where the similarity to the target distribution will usually be the utility (e.g., similarity to Wikipedia data). Other distribution matching methods use the statistics of data sampled from the target dataset as the utility (e.g., total number of characters per example).
Distribution diversification methods (Figure 2, right) aim to prioritize heterogeneity in a sample, removing redundancies. They operate within a representation space that allows for similarity to be measured across data points. The utility of a data point is defined by its relation (similarity) to the other data points in the representation space. Distribution diversification methods often remove data points that are similar in some representation space (e.g., characters or vectors). By removing data points that are very similar, diversification methods can remove redundancies in the data, leading to improved training efficiency by reducing the dataset size. Additionally, because distribution diversification methods force the dataset distribution assign probability to a broader range of datapoints, they can lead to decreased memorization, decreased bias, and improved robustness.
@@
Methods that alter the dataset aim to increase or decrease the frequency of individual data points within the dataset in order to increase (decrease) the resultant distribution density around desirable (undesirable) regions. These methods take as input a dataset and assign each individual data point a non-negative integer value representing the number of times that data point should be included in the resultant dataset. Formally, a dataset distribution altering function $\phi_{j}:\mathcal{D}^{N}\to\mathbb{N}_{0}^{N}$ maps each data point $x^{(i)}\in\mathcal{D}$ to a non-negative integer based on the utility function, $\upsilon(x^{(i)})$. Each datapoint is then filtered ($\phi_{j}(x^{(i)})=0$), maintained ($\phi_{j}(x^{(i)})=1$), or oversampled ($\phi_{j}(x^{(i)})>0$) according to the utility function.
Methods that alter the data point aim to adjust the content within a data point to better match a desirable distribution of tokens so that the individual data points within a dataset more closely resemble the desired characteristics. In practice, this can take the shape of removing individual lines, or chunks of text, from a document. For example, when training a model for natural language, it can be beneficial to remove HTML tags as they are likely to be out of distribution if the desired target domain is natural language.
@@
Methods that alter the dataset distribution lead to another dimension of variance: whether the selection function $\phi$ assigns binary values (i.e., include or remove) or whether larger integer values are permitted. Methods that only assign binary values are generally referred to as filtering, and the goal is usually to adjust the dataset distribution by removing undesirable data points from the dataset. On the other hand, methods that can assign any natural number are commonly referred to as data mixing, and their goal is often to adjust the dataset distribution by prioritizing certain subsets of the dataset that have high utility values, while decreasing the distribution density around subsets with lower utility. It is very common for data mixing methods to define a utility function that assigns the same utility to entire subsets of data (e.g., all web text gets the same utility, and all books get a different utility). Filtering and mixing can be used as part of a pipeline, where data is first filtered and then mixed.
2.3.4 Training Stage
Each training stage has different goals, and so the data selection methods for each stage will use different mechanisms to achieve those goals. For example, the target distribution is fairly well-defined when training a language model for a specific domain or task, but much less defined when pretraining a general purpose language model, thus data selection methods can utilize this information to target better data. Another consideration is the number of candidate data points at each training stage. For example, pretraining likely has a significantly larger number of candidate data points than instruction-tuning, motivating the need for efficient data selection methods in pretraining, while instruction-tuning can afford more expensive methods.
Data selection has been studied, to some extent, in all stages of training. However, information on data selection for pretraining is most limited, and therefore we dedicate the main portion of this work to comparing, contrasting, and better understanding the methods of data selection in pretraining, with more concise sections discussing data selection for other training stages.
3 Data Selection for Pretraining
@@
The goal of pretraining is usually to train a “general-purpose” model, which requires training on massive quantities of text (often trillions of tokens for recent large language models). Selecting the best data from such large quantities can be very expensive, so a common first step in the process is to remove data with various filters, and it is common to apply multiple filters that are pipelined together to achieve the desired dataset. The order in which we present pretraining data selection methods is, roughly, based on the order that they are used within real data selection pipelines. Of course, not all pipelines require every method presented here, and depending on the case the exact ordering may differ slightly. We provide a high level overview of the filtering pipline in Figure 4, and references to relevant subsections as well as axes of variation in Table 1. The pretaining data selection step illustration is depicted in Figure 3.